OpenStudy (anonymous):

A seconds pendulum is a pendulum whose period is precisely two seconds; one second for a swing in one direction and one second for the return swing. The length of a seconds pendulum in Tokyo, Japan is 0.9927 m and that in Cambridge, England is 0.9942 m. The ratio of the acceleration due to gravity at the two places, gT /gL is

OpenStudy (anonymous):

0.9987. 1.0150. 1.0008. 1.0015. 1.0000. which one is correct !!

OpenStudy (anonymous):

The period of the pendulum is given by: \[\huge{T=2\pi \sqrt{\frac{L}{g}}}\] Since the period of each pendulum is the same you can set them equal to each other through equations: \[\huge{T_{T}=T_{C}}\] \[\huge{2\pi \sqrt{\frac{L_T}{g_T}}=2\pi \sqrt{\frac{L_C}{g_C}}}\] From here you should be able to rearrange to solve the ratio: \[\huge{\frac{g_T}{g_C}=?}\]

OpenStudy (anonymous):

is the period of a pendulum a given formula?

OpenStudy (anonymous):

yes...see first equation above. T = Period L = Length g = Force of gravity

OpenStudy (anonymous):

so you will have to memorise it

OpenStudy (anonymous):

I'm not sure what you'll have to do. It depends on the context in which you were taught it i suppose